OpenStudy (anonymous):

please help will medal True or false The domain of y=cot x is given by x=/ npi

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@wio

OpenStudy (dtan5457):

@Nnesha

Nnesha (nnesha):

can you wait i gtg this time :(

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

The domain, is all the possible independent variable values, in this case the x values that will result in an output y value

OpenStudy (danjs):

\[\cot(x) = \frac{ \cos(x) }{ \sin(x) }\]

OpenStudy (danjs):

When the sin(X) = 0, the cotangent function will be undefined

OpenStudy (danjs):

sin(x) = 0 for x = pi, 2pi,3pi,...,n*pi

OpenStudy (danjs):

so the cot(x) will be undefined at all values of x, where x is some multiple of pi.

OpenStudy (danjs):

therefore the domain of Cot(x) is, All x, such that x does not equal n*pi, where n=0,1,2,3....

OpenStudy (danjs):

so true

OpenStudy (anonymous):

Ok thanks I was pretty sure it was true but i wanted to see if I was right

OpenStudy (danjs):

GRAPH of Y= SIN(x) Y=SIN(x) = 0 , for X= multiples of pi